OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)) y=(2/x^2) + 3

OpenStudy (anonymous):

there are lots of possible answers

OpenStudy (anonymous):

i will give two: \[g(x)=x^2, f(x)=\frac{2}{x}+3\] \[g(x)=\frac{2}{x^2}, f(x)=x+3\]

OpenStudy (anonymous):

i am sure there are more

OpenStudy (anonymous):

@satellite73 thank you! I was thinking about pulling out 'x^2' to be g(x) but I wasn't sure if I was doing it right or not.

OpenStudy (anonymous):

yw